### v1

# A Comprehensive Benchmarking Analysis of Fault Recovery in Stream Processing Frameworks

Abstract.
Nowadays, several software systems rely on stream processing architectures to deliver scalable performance and handle large volumes of data in near real time. Stream processing frameworks facilitate scalable computing by distributing the application’s execution across multiple machines. Despite performance being extensively studied, the measurement of fault tolerance—a key and most appealing feature offered by stream processing frameworks—has still not been measured properly with updated and comprehensive testbeds. Moreover, the impact that fault recovery can have on performance is mostly ignored. This paper provides a comprehensive analysis of fault recovery performance, stability, and recovery time in a cloud-native environment with modern open-source frameworks, namely Flink, Kafka Streams, and Spark Structured Streaming. Our benchmarking analysis is inspired by chaos engineering to inject failures. Generally, our results indicate that much has changed compared to previous studies on fault recovery in distributed stream processing. In particular, the results indicate that Flink can be the fastest and stablest under failures. Moreover, Kafka Streams shows performance instabilities after failures, which is due to its current repartitioning strategy that can be suboptimal in terms of load balancing. Spark Structured Streaming shows suitable fault recovery performance and stability, but with higher event latency. Our study intends to (i) help industry practitioners in choosing the most suitable stream processing framework for efficient and reliable executions of data-intensive applications; (ii) support researchers in applying and extending our research method as well as our benchmark; (iii) identify, prevent, and assist in solving potential issues in production deployments.
benchmarking, fault tolerance, performance, stream processing
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software performanceccs: Computer systems organization Cloud computingccs: Information systems Stream management
1. Introduction
Stream processing refers to a computing paradigm and architecture pattern for building distributed, event-based software systems that process massive amounts of data in near real time (Hirzel et al., 2014). Today’s stream processing paradigms enable software engineers to model their applications’ business logic as high-level representations in a directed acyclic graph (DAG), without explicitly defining the physical execution plan. This high level of abstraction is provided by industry-grade, open-source stream processing frameworks such as Spark (Armbrust et al., 2018), Flink (Carbone et al., 2015), or Kafka Streams (Wang et al., 2021). Such frameworks support software engineers in building highly scalable and efficient applications that process continuous data streams of massive volume. They provide high-level APIs and domain-specific languages to define the processing logic as directed acyclic processing graphs that filter, transform, aggregate, and merge data streams.
Over the last decade, several works have been published that evaluate the performance of distributed stream processing frameworks (Vogel et al., 2023; Henning and Hasselbring, 2024; Henning et al., 2024) or optimizations (Hirzel et al., 2014; Herodotou et al., 2020; Vogel et al., 2022). Beyond measuring performance metrics such as throughput and latency, fault tolerance stands as a critical requirement for continuously-operating production systems. Failures can occur unpredictably across various levels, from physical infrastructure to software layers. Stream processing systems, designed for continuous, low-latency processing, demand swift recovery mechanisms to tolerate and mitigate failures effectively.
It is notable that a relevant number of studies reported fault tolerance as challenging for stream processing (Vogel et al., 2023). Although fault tolerance guarantees such as at-least-once or exactly-once are already provided by modern stream processing frameworks (Fragkoulis et al., 2024) and partially evaluated in the literature (van Dongen and van den Poel, 2021; Wang et al., 2022), there are still many open questions for stream processing in scalable industry-grade deployments: i) With the new developments of fault tolerance features, how fast do the executions recover from failures? ii) What is the failures’ impact on the application performance? iii) What is the impact of more critical failures such as correlated failures of many worker instances (entities executing the stream processing applications)? iv) How stable is the fault recovery across recurrent failures (we call it stability)?
Considering the open questions of stream processing frameworks’ resilience to failures, we aim to extend the current methods for evaluating fault recovery and provide updated insights from conducting comprehensive experimental analysis. Particularly, this paper provides the following contributions to industry practitioners and the research community:
  • A comprehensive methodology for measuring fault recovery. The methodology covers stream processing’s most relevant metrics (Vogel et al., 2023), new metrics to improve fault recovery evaluations, a method for failure injection, and a method for automating fault recovery time measurements.
  • An extension of ShuffleBench (Henning et al., 2024) for fault tolerance measurements in the different state-of-the-art stream processing frameworks, tooling to automate benchmarks’ execution, and data collection in Kubernetes-based cloud environments.
  • An experimental analysis covering application metrics, system resources utilization, and fault recovery. [^1] On publication of this study, we will also provide as supplemental material an extended replication package of ShuffleBench for fault tolerance measurements. Moreover, the experimental data is also to be shared such that other researchers and practitioners may repeat and extend our work. Our evaluation covers the open-source stream processing frameworks Flink, Kafka Streams, and Spark Structured Streaming due to their industry acceptance and academic relevance (Vogel et al., 2023).
2. Background
This section briefly introduces the fundamental concepts of modern stream processing frameworks and fault tolerance.
2.1. Distributed Stream Processing
Stream processing frameworks perform operations such as filterings, transformations, or aggregations in near real-time on continuous streams of data (Hirzel et al., 2014). State-of-the-art frameworks are designed for high throughput and low-latency processing, while also scaling with massive amounts of data (Fragkoulis et al., 2024; Henning and Hasselbring, 2024). To address these requirements, they run in a distributed fashion on commodity hardware. A key advantage of stream processing frameworks is that they provide dataflow models that abstract aspects such as cluster management, state management, and time semantics from their users (Sax et al., 2018). With such models, engineers describe the processing logic in DAGs of processing operators. The frameworks allow the initiation of multiple worker instances across various compute nodes with multiple threads, each instance handling a distinct portion of the data.
While the isolated processing of data records remains unaffected by the assignment of data portions to worker instances, processing that depends on previous data records, such as aggregations, requires state management. Similar to the MapReduce (Dean and Ghemawat, 2008) programming model, keys are assigned to records before a stateful operation. This allows the stream processing frameworks to route all records with the same key to the same instance, where state synchronization among instances can be avoided. When a processing operator modifies the key of a record and a subsequent operator performs a stateful operation, the framework divides the dataflow graph into subgraphs that can be independently processed by different worker instances. Popular stream processing frameworks include Apache Flink (Carbone et al., 2015), Apache Kafka Streams (Wang et al., 2021), and Apache Spark with its Structured Streaming engine (Armbrust et al., 2018).
2.2. Fault Tolerance
Achieving consistency in stream processing has posed a longstanding research challenge, which is partly attributed to the lack of a formal problem specification (Wang et al., 2021). Although executions can be subject to many failures, we expect computing systems to produce correct results despite the failures. Such a requirement instigated research on fault tolerance, where notable advances achieved in distributed systems are applied to stream processing (Wang et al., 2021).
In the absence of fault tolerance, stream processing applications would be forced to restart data processing from scratch whenever there is a loss due to a failure. Given that applications may have varying fault tolerance needs, the stream processing domain focuses on different processing semantics that precisely describe how a system is affected by failures (Fragkoulis et al., 2024). The relevant terms to characterize the fault tolerance semantics guarantees are at-most-once semantics (AMOS), at-least-once semantics (ALOS), and exactly-once semantics (EOS). Currently, the majority of stream processing applications demand reliable executions with strong semantics guarantees. Thus, AMOS is obsolete (Fragkoulis et al., 2024) and we focus on ALOS and EOS.
At-least-once semantics
ALOS ensures that the system maintains consistent results even in the event of failures, preventing data loss. However, during recovery, duplicate output may occur as records might be processed more than once (Fragkoulis et al., 2024). The frequency of such duplicates largely depends on the system’s implementation, the type of failure encountered, and the timing of the failure during execution. Additionally, handling potential duplicates can be accomplished externally, for instance, through downstream systems filtering duplicates or database overwrites. With effective implementations, the fault tolerance guarantee provided by ALOS can be comparable to EOS.
Exactly-once semantics
EOS refers to the ability of the system to process the records exactly once without lost or duplicated results even under failures. Therefore, EOS requires additional implementation and has a performance impact on records processing (van Dongen and van den Poel, 2021).
In this paper, we focus on semantics guarantees regarding the system output because this is the most relevant for the majority of use cases and also covers the guarantees in the internal state updates. The stream processing frameworks provide different recovery semantic guarantees. This is important because the required guarantee varies according to the stream processing application and its use cases. In some scenarios, guaranteeing exactly-once can be of paramount importance, e.g., avoiding duplicate alerts in an observability platform. In other cases, at-least-once with rarely duplicated outputs is not equally detrimental, e.g. a monitoring system that receives twice the same CPU metric of a timestamp. We intend to measure the impacts of the stream processing frameworks when running with the fault recovery guarantees. Importantly, correct fault recovery is a crucial component of fault tolerance. While fault recovery focuses on restoring normal operation after a failure, fault tolerance aims to prevent downtime for high availability.
3. Research Landscape
This section overviews the related literature and we discuss existing research and analysis gaps.
3.1. State-of-the-art
Recent publications have presented various experimental evaluations of fault recovery in stream processing (Vogel et al., 2023). Although many studies consider failures of worker instances (Lu et al., 2014; Qian et al., 2016; Lopez et al., 2016), the most updated and relevant is (van Dongen and van den Poel, 2021). It dives into the implementation, performance, and efficiency of fault recovery in four state-of-the-art stream processing frameworks: Spark Streaming, Flink, Spark Structured Streaming, and Kafka Streams. In addition, they evaluated the behavior of these frameworks under different types of faults and settings, including master failure with and without high-availability setups, driver failures for Spark frameworks, worker failure with or without exactly-once semantics, and application failures.
Wang et al. (2022) propose a taxonomy of fault tolerance in stream processing. They also propose an evaluation framework tailored for fault tolerance, demonstrating experimental results on two representative open-source stream processing frameworks. Potential limitations in the current approaches are also discussed.
Theodorakis et al. (2021) discuss the challenges and solutions in making single-node stream processing fault-tolerant while maintaining high performance. Due to the limited I/O bandwidth of a single node, it becomes infeasible to persist all stream data and operator state during execution. To address this issue, Scabbard was proposed as the first single-node framework that supports exactly-once fault tolerance. However, the scalability of Scabbard can still be somewhat limited to a single node.
Wang et al. (2021) explain Kafka Streams’ processing logic and demonstrate how Kafka Streams behaves in practice with large-scale deployments and performance insights exhibiting its flexible and low-overhead trade-offs. A very recent paper from Siachamis et al. (2024) focuses on complementary aspects of fault tolerance. The authors explored three approaches for checkpoint protocols (coordinated, uncoordinated, and communication-induced) aimed at fault tolerance in stream processing. The analysis delved into the theoretical strengths and weaknesses inherent in each approach. Their findings not only validated the prevailing preference for the coordinated approach but also illustrated instances, where the uncoordinated approach exhibited resilience and better performance.
Considering that many optimizations can be done for fault tolerance (Hirzel et al., 2014; Vogel et al., 2022), Su and Zhou (2021) address the issue of recovering from correlated failures in large-scale clusters. The authors propose an incremental and query-centric recovery paradigm. They formulate the problem of recovery scheduling under correlated failures and design algorithms to optimize the recovery latency.
3.2. Research Gaps
While fault recovery in stream processing has been addressed in various related studies, our understanding is that the methods and insights provided offer limited coverage. Despite being recognized as essential within a leading cloud observability platform, we identified significant research and analysis gaps in the state-of-the-art regarding fault recovery (Vogel et al., 2023):
  • Considering the continuous development of fault tolerance features and optimizations by the frameworks’ open-source communities, there is a need for updated measurements on frameworks’ executions facing worker instances’ failures.
  • The few studies that compare the frameworks regarding worker instances’ failures cover a single failure (van Dongen and van den Poel, 2021; Wang et al., 2022). However, a large physical machine crash impacts several smaller virtual machines or pods running within it. What can happen is that more than one worker instance fails. However, it remains unclear what is the impact of such a more critical failure in the running stream processing applications.
  • The existing literature only considers that worker instances can fail a single time (van Dongen and van den Poel, 2021; Wang et al., 2022). In real-world industry deployments, failures happen recurrently and at unpredictable times. One can assume that the frameworks’ recovery will always be the same, but we observed that such an assumption can have weak coverage with the reality of large deployments. This occurs mostly because the impact that a given failure has in the applications varies according to the exact moment that the failure happens (e.g., how far from the last saved state checkpoint), what working instances are affected by the failure, and the effectiveness of the subsequent recovery (Geldenhuys et al., 2022). Therefore, the level of fault recovery stability across recurrent failures remains unclear.
4. Experimental Setup
LABEL:{sec:gaps} described existing research gaps. Here we show our experimental setup to extend the fault recovery measurements.
4.1. Benchmark
ShuffleBench (Henning et al., 2024) is a benchmark for evaluating the performance of modern stream processing frameworks. It focuses on use cases where stream processing frameworks are mainly employed for shuffling (i.e., re-distributing) data records to perform state-local aggregations, while the actual aggregation logic is considered as a black-box software component. ShuffleBench is inspired by requirements for near real-time analytics of a large cloud observability platform, but it is highly configurable to allow domain-independent evaluations. It comes as a ready-to-use open-source software [^2] https://github.com/dynatrace-research/ShuffleBench utilizing existing Kubernetes tooling and providing implementations for different state-of-the-art stream processing frameworks.
ShuffleBench utilizes and extends the benchmarking framework Theodolite (Henning and Hasselbring, 2022) to automate the benchmark execution in Kubernetes-based cloud environments. Figure 1 depicts our dataflow architecture for a corresponding stream processing application. In short, data records are read from a messaging system (Kafka in our case), assigned to keys, and shuffled such that all records having the same key are forwarded to the same worker instance. There, a key-specific custom aggregation is performed and, optionally, an output event is created that is then written back to Kafka. We refer to our previous work for further details about ShuffleBench (Henning et al., 2024).
Figure 1. The ShuffleBench dataflow architecture of a stream processing application with three worker instances (Henning et al., 2024).
ShuffleBench has a flexible architecture and supports several parametric configurations. In this work, we extend ShuffleBench for fault tolerance measurements. Figure 2 provides an overview of ShuffleBench’s benchmark components and their extension to fault tolerance, where the scenarios and metrics are described in Section 5. We conduct our experimental evaluation in a Kubernetes cluster managed by the Elastic Kubernetes Service of Amazon cloud with the following pods running in cloud instances: 2 load generators (m6i.xlarge instances), 2 Kafka brokers (m6i.2xlarge), a manager node used on Flink and Spark (m6i.large), and 8 pods for worker instances with 1 CPU, 3 GB of RAM (running on c5.large instances).
In our experiments, we set up 20 000 real-time consumers that all have the same selectivity, which sum up to 50 %, meaning that each record is forwarded to $0.5$ consumers on average. Each consumer emits an output event for every record received. To consume and output results, 40 input and output Kafka partitions were used to maintain a suitable parallelism level that minimizes overheads. We run each experiment for 56 minutes in such a way that the benchmarks run for long enough to extract meaningful quantitative data, test with recurrent failures injected, and have sufficient time for every subsequent recovery.
Figure 2. Overview of ShuffleBench benchmark components. Green components are extensions to our previous work (Henning et al., 2024).
4.2. Frameworks Deployment
We configured the frameworks on ShuffleBench deployment and parameters described in Section 4.1. To support fault tolerance, Flink (v. 1.17) and Spark Structured Streaming (v. 3.4) had checkpointing enabled for at-least-once semantics (ALOS), see more in LABEL:{sec:background-fault-tolerance}. Kafka Streams (v. 3.5) by default supports ALOS. Fault tolerance on Flink and Spark Structured Streaming involves periodic state backup via checkpointing, demanding a reliable filesystem to store state copies (van Dongen and van den Poel, 2021). We used in our setup an Amazon Elastic File System (EFS) filesystem volume mounted in the pods.
Kafka Streams provides fault recovery with its strong integration with Kafka. Kafka acts as a messaging layer to safeguard state information. In-memory state stores are associated with a replicated changelog topic within Kafka, which tracks state updates. Whenever a task interacts with the store, it maintains a changelog. If a task encounters failure, the system automatically restarts it on one of the remaining worker instances. The state stores can be restored by replaying their changelog, ensuring data consistency. Kafka Streams provides a configuration to enable exactly-once (exactly_once_v2), which was enabled to evaluate exactly-once. Exactly-once in Flink can be enabled in a similar way by enabling the DeliveryGuarantee configuration to exactly_once.
Flink’s checkpointing interval and Kafka Streams’ commit interval were set to 2 seconds. Although such value can be tuned (Geldenhuys et al., 2022), it is comparable to Spark executions’ micro-batching interval after the warmup in our setup. This is relevant for having executions more comparable because Spark Structured Streaming saves a checkpoint when the batch is completed. Flink was also configured to run in the reactive mode, which is a modern feature necessary for supporting all the failures injected in our experiments.
Write-ahead-logs (WAL) was enabled in Spark Structured Streaming. Moreover, the number of SQL shuffle partitions was set to 16 to balance between the parallelism level and the resources. The Spark manager pod needed additional containers, the spark driver (execution manager), and Kafka offset committer that we implemented to support application metrics measurements in a comparable way to the other frameworks.
5. Experimental Method
This section describes our proposed method to evaluate fault recovery on stream processing.
5.1. Goals
Considering the best practices and the existing research gaps described in Section 3.2, the following are the goals of our research method: i) Create comparable deployments across different frameworks with similar throughput processing capabilities; ii) Measure with comprehensive evaluation how the frameworks react to failures, covering periodic and recurrent failures of different types; iii) Have evaluations that are realistic to real-world scenarios, focusing on the most likely failures affecting worker instances.
5.2. Failure Injection Scenarios
Experimental scenarios were inspired by chaos engineering (Basiri et al., 2016). Failures are injected using Chaos Mesh, [^3] https://github.com/chaos-mesh/chaos-mesh an open-source chaos engineering platform integrated with the Kubernetes deployment. We consider two scenarios:
Pods kill:
Random pods executing worker instances are deleted. When the pods are no longer accessible, both the records they were processing and their associated state are lost. Therefore, the frameworks have to trigger recovery from such a failure. Deployed with Kubernetes, new pods will be created and added to the pool of worker instances within an interval of seconds (usually taking from 2 to 10 seconds). We experimented with killing 1, 2, and 4 worker pods at once. Such values were determined to simulate simple and critical failure scenarios. Such failures are recurrently injected every 5 minutes to evaluate the fault recovery stability. This interval was defined for a reasonable recovery time between failures and also to avoid unnecessary long executions that would waste resources.
Pods crash:
Random worker instances pods face an operating system crash. This is expected to further test the resilience because the availability of resources is impacted. Contrasting with pods kill, in the pods crash failure no new pods are created to replace the failed ones. This failure also is expected to cause loss in the internal state of the worker instances affected, demanding failure recovery. Every failure impacts two random worker pods, which is a value set to represent a critical failure. The failures are recurrently injected every 10 minutes to evaluate the fault recovery stability and the failure remains for 5 minutes (value set according to feedback of industry partners). Then, executions can resume the normal state. Moreover, 5 minutes is a suitable value for creating and adding new virtual machines (Su and Zhou, 2021) to the pool.
5.3. Metrics
We collected several metrics of the benchmarking application. The metric to measure processing speed is throughput, measured in records processed per time unit and collected every second as a moving average of the last 5 seconds. A higher throughput is usually desired in many use cases. We measure input throughput refers to consuming from the Kafka input topic, and output throughput which is the application output written into another Kafka topic. We also measure event latency (the lower latency tends to be better) as an average of a 10-second window. We show latency percentiles to measure the distribution of events times: the p50 (50th percentile), p90, and p99. We also measure the consumer lag corresponding to the number of records buffered in the input topic.
The fault recovery on stream processing is measured with the metrics: Fault recovery performance, the impact of fault tolerance guarantees, stability, and fault recovery time (see Figure 2). We also present the following system monitoring indicators collected every 2 seconds from the cloud instances that run the worker pods: CPU utilization, memory utilization, and network traffic. [^4] To enhance visual clarity, the CPU utilization and network traffic of worker nodes are displayed using a 10-value moving window average.
6. Fault Recovery Characterization
This section shows results demonstrating how the injected failures impact the frameworks’ execution as well as how they react to the failures. We provide results on the application metrics and system resource utilization (see more in Section 5.3). This section intends to demonstrate the potential impacts of faults and their recovery on the different stream processing frameworks and their Quality of Service (QoS). In this section, we show first failures for the sake of visual clarity, Section 7 extends these characterizations providing results with up to 10 recurrent failures injected.
6.1. ALOS Recovery Characterization
(a) Fault recovery
(b) Resources utilization
Figure 3. ALOS fault recovery characterization: Random pods executing worker instances are killed, repeated 3 times.
Figure 3 shows how the executions with the different frameworks behave when pods executing worker instances are killed (represented by the red line). To have more comprehensive characterizations, we present results with three failures that occur periodically with an interval of 5 minutes. It is also important to note that the plots show different metrics (see y-axis labels) over time.
When the failures occur, the worker instance impacted is lost. Then, the Kubernetes infrastructure creates a new pod that is added to the pool of worker instances. Moreover, the stream processing frameworks are expected to start scheduling computations to new worker instances. This is achieved in Flink and Kafka Streams by rebalancing the partitions to worker instances. In Spark Structured Streaming, the worker instances are added to the pool of executors and they start receiving tasks to compute.
Fault recovery observation 1: The frameworks recover correctly from the failures. From a quality of services perspective, it is noticeable that Flink executions recover quicker from failures in comparison to Kafka Streams and Spark Structured Streaming, which is evinced by a shorter downtime in the output throughput and quicker reduction of the latencies.
Performance observation 1: After failures, Flink can present fluctuations in the data consumed from Kafka. To the best of our knowledge, this phenomenon has not been previously reported. The variation in the amount of data ingested after failures may be attributed to the uncertainty of the timing of failures relative to the previously saved checkpoints. However, we have deferred further exploration of this phenomenon to future work, as it does not significantly impact the output throughput or the latency.
Performance observation 2: Kafka Streams achieves a stable throughput with and without failures of one worker instance pod, but after failure has fluctuations in resources usage. Kafka Streams is the one that accumulates more lag after failures, which is mostly due to its default partition assignment strategy triggered after the failures. After the first failure, there are fluctuations in the CPU utilization and network traffic in the worker machines, which indicates a potential imbalance. This also relates to the partition assignment strategy where some worker instances in the consumer group are assigned to more partitions or partitions with more lag accumulated. See more in the resource consumption observation 2.
Performance observation 3: Spark Structured Streaming’s dynamic definition of micro-batch intervals is finding suitable configurations. Dynamic micro-batching is evident in the first few minutes of the execution with the throughput and lag fluctuating significantly. Such fluctuations happen because the micro-batch interval is too large causing more data to be consumed and batches being triggered less frequently. After a few minutes, the micro-batch interval is self-regulated to close to 2 seconds, which is a configuration suitable to sustain the input throughput and maintain a reasonable latency (the p50 latency is comparable to the micro-batch interval). Still, Spark’s latency is significantly higher than the other frameworks that process the records based on events instead of micro-batches. This outcome is consistent with the related literature (van Dongen and van den Poel, 2021).
(a) Fault recovery
(b) Resources utilization
Figure 4. ALOS recovery characterization: Two random pods executing worker instances are killed, repeated 3 times.
Resource consumption observation 1: Spark Structured Streaming has the most dynamic CPUs utilization. This is due to its modern execution model that processes based on fine-grain tasks assigned to the spark’s executors (worker instances). Tasks are created within the so-called stages that execute the micro-batches, the number of such tasks is determined based on the number of partitions (consumed from Kafka and subsequent shuffle partitions between DAG’s stages). Such a dynamic execution model can be more scalable and suitable for effective load-balancing but at the cost of potential overheads of having to assign all the fine-grain tasks to the executors. It is also noticeable that, from the operating system view, Spark Structured Streaming execution is the one that allocates the highest memory space. [^5] This memory observation should not be taken as evidence because the way the Java virtual machine allocates memory and computes can vary across the frameworks. Therefore, the operating system view of the memory usage does not guarantee accuracy. Considering that memory consumption can have a heavy impact on performance and cost, this motivates future analysis to comprehend the root causes and potential mitigations. One direction is to measure the garbage collection (GC) effectiveness and performance, where the Spark’s GC was fine-tuned in related works (van Dongen and van den Poel, 2021). Moreover, this behavior could be a potential reason for Spark’s performance limits found in our previous work (Henning et al., 2024).
Resource consumption observation 2: From a resource consumption perspective, Kafka Streams shows unstable behavior after the failures are injected. This is supported mostly by varied CPU utilization and network traffic in the worker nodes, which suggests some imbalance that we believe to be load-related when repartitioning after the failures. This outcome aligns with performance observation 2: the partition assignment strategy can be suboptimal after failures making some workers overloaded and others underutilized. Furthermore, from an operating system view, Kafka Streams execution is the one allocating least memory, which is a positive outcome from a resource efficiency view.
Resource consumption observation 3: Flink shows consistent resource utilization under failures. There are variations caused by the failures, but the execution returns to a normal stable state. It is worth noting that Flink’s execution is the one causing more CPU utilization in the Kafka Broker servers, which are already large-sized instances. This may be attributed to Flink’s fine-grained event-based processing, which continuously generates outputs to the Kafka topic. We noted that Flink’s performance may rely more than other frameworks on the behavior of the Kafka brokers and expect to address this in a future analysis.
Figure 4 shows the scenario where periodically two random pods executing worker instances are killed. We expect this scenario to be more challenging as more workers failing means that more internal state can be lost and more records need to be replayed from the input source and reprocessed. Compared to Fig. 3, Fig. 4 shows similar performance, fault recovery, and resource consumption observations. However, it is noticeable in Fig. 4 that as Kafka Streams’ execution accumulates more lag after the failures, its event latency increases dramatically.
(a) Fault recovery
(b) Resources utilization
Figure 5. ALOS fault recovery characterization: Two random pods executing worker instances crash for 5 minutes.
With a more critical failure as the one with two pods failing, the unbalances caused by Kafka Stream’s repartitions after failures evinced in Performance observation 2 and Resource consumption observation 2 become more critical and impactful. On the one hand, Kafka Streams strategy can evenly distribute the partitions to the available workers in some cases of repartitions after failures. On the other hand, there are many cases where the distribution of partitions is suboptimal. For instance, the results from Fig. 4 indicate that after the first failure, the repartition was suboptimal with some workers taking too long to process the lag accumulated. However, after the third failure, the repartition seems to be optimal in such a way that the lag is fast processed and the latency reduced. This is supported by the resource utilization data that shows that a worker node (green color) consumes more CPUs and has higher network traffic while the lag is being reduced. Potentially this specific node was running a pod of a worker instance assigned to a lagged partition that was able to process the lagged records. Therefore, Fault recovery observation 2 is that Kafka Streams’ fault recovery can have an inconsistent impact on the applications’ performance across recurrent failures.
Although four worker instances failure is a more critical scenario, the results are aligned with the insights from the failures of two worker instances. Therefore, for the sake of conciseness, we do not show here results with 4 pods being killed.
Figure 5 shows the scenario where the failure injected is that random pods executing worker instances face an operating system crash (see Section 5.2 for further details and motivations for this scenario). As they become unavailable, the records being processed by them and their local state is lost. In contrast with the pods killed failure, in this case, no new pods are created. Therefore, the execution has to continue processing with fewer worker instances.
Focusing on Fig. 4(a) after the failure is injected, in Fig. 4(b) significant drops in system utilization are noticeable in the machines executing the two crashed pods. On the other hand, the active worker instances in all three frameworks increase their resource utilization to cope with the application processing demand. When the failure injected is halted (the gray line), the executions can return to the normal state with crashed worked instances being re-added to the workers pool. From the performance and QoS standpoint evinced in Fig. 4(a), Fault recovery observation 3 is that the executions of the three frameworks recovered from a critical failure that reduces the resource availability, and when the failure is halted they returned to a comparable state as the one before the failure.
One can note that Flink’s execution is the one that recovers faster to a normal performance state. Kafka Streams is the one accumulating more lag and its execution took almost 3 minutes to normalize the latency. Spark Structured Streaming demonstrates fluctuations after the failure, which is due to reprocessing of the tasks that were assigned to the failed executors and due to the dynamic micro-batching. With fewer executors and consequently fewer worker instances resources, Spark Structured Streaming stabilized the micro-batching interval with a higher value of around 3 seconds. The increase in the micro-batching interval was seemingly necessary to cope with the input load speed. However, this results in a slightly higher latency. Moreover, it is noticeable that Spark Structured Streaming’s output throughput faces more fluctuations after the failure is injected.
After the failure is halted the executions with the three frameworks return to normal states. Comparing the different worker instances, Kafka Streams ends up with some noticeable imbalances in network traffic and CPU utilization (Fig. 4(b)).
6.2. EOS Recovery Characterization
Figure 6. EOS recovery characterization: Random pods executing worker instances are killed, repeated 3 times every 5 minutes.
Supporting exactly-once semantics (EOS) demands further configurations and deployments. One more Kafka broker was necessary for resource stability and the fault tolerance guarantee had to be configured in the frameworks. Moreover, as also done in the related literature (van Dongen and van den Poel, 2021; Wang et al., 2021) when testing EOS, the configuration of the isolation level is set to read only committed data, which prevents the reading of duplicated messages in downstream operators.
Considering that Spark Structured Streaming can not natively guarantee exactly-once when writing the output to a Kafka topic, we adopt the same strategy of the related literature (van Dongen and van den Poel, 2021) and do not present Spark Structured Streaming with exactly-once.
For conciseness, we show only the results with one pod kill that already provides the insights necessary to highlight Flink’s advantages and Kafka Streams’ shortcomings when EOS is needed. In Fig. 6 it becomes evident that Flink, compared to Kafka Streams, provides better performance and fault recovery, [^6] We do not show the results from machine system monitoring of resources because it becomes repetitive and without new insights w.r.t. the previous results. thus resulting in Fault recovery observation 4: Flink provides better performance and fault recovery for use cases that demand exactly-once guarantees.
Focusing on the output throughput, both frameworks’ executions are similar until the first failure occurs. Then, Flink recovers faster in less than a minute to a normal output throughput that becomes stable. On the other hand, Kafka Streams took more than one minute to reach a necessarily higher output throughput. Moreover, after the failure, Kafka Streams’ output throughput is unstable and subject to fluctuations. We believe that this is due to the lag accumulated and being too slowly processed. This aligns with Performance observation 2 and Fault recovery observation 2. However, the scenario with exactly once is a more critical one, which results in a stronger impact of Kafka Streams’ suboptimal implementations.
One can note that Flink achieved a much lower latency in Fig. 6. Kafka Streams stabilized in a latency above 2 seconds and after the first failure, the latency does not normalize due to the large amount of lagged records. After a failure, Kafka Streams’ latency exceeds the 10-second metric bucket limit (Henning et al., 2024), rendering it invisible in the y-range plotted. These observations in Kafka Streams are, to the best of our knowledge, novel findings w.r.t. related approaches covering Kafka Streams with EOS (van Dongen and van den Poel, 2021; Wang et al., 2021). This can be mostly due to the fact the Kafka Streams was only tested under failures in (van Dongen and van den Poel, 2021) with a single failure injected. Moreover, technical reasons for the higher latency are due to the records buffering in Kafka Streams’ repartition topic and that its partition assignment strategy has been changed since (van Dongen and van den Poel, 2021), which implies that the current strategy is suboptimal for rebalancing after failures.
7. Fault Recovery Stability
This section presents the results of fault recovery stability under recurrent failures injected. Note the different ranges and scales on the y-axis. For conciseness, we limit the discussion to the results with ALOS guarantees as those are the most complete and insightful.
(a) Flink
(b) Kafka Streams
(c) Spark Structured Streaming
Figure 7. ALOS fault recovery: Random pods executing worker instances are killed, repeated 10 times every 5 minutes.
Figure 7 complements the results from Fig. 3 with a focus on the fault recovery stability. Fig. 6(a) shows that despite some fluctuations in the input throughput of Flink’s execution, it achieved a stable behavior under recurrent failures. Figure 6(b) shows Kafka Streams execution in a scenario where one random worker instance is recurrently killed. Focusing on throughput and lag, it becomes evident that Kafka Streams failure recovery follows a pattern with different phases after the failures and it lasts for several minutes.
Although Kafka Streams’ failure recovery pattern is consistent across the different failures, it varies significantly for how long the patterns and their phases occur, and their impact on performance (e.g., lag peaks and throughput drops). Moreover, the fluctuations in lag and throughput have a detrimental impact on the latencies, evincing limitations in Kafka Streams’ fault recovery. On the one hand, in the recovery of the sixth failure shown in 6(b), the lag accumulated as well as the throughput are lower, and the latency increase is not disruptive with the p50 being kept within the boundaries. On the other hand, following the seventh failure, the performance degradation becomes critical. Consequently, the execution takes around 4 minutes to return to a state comparable to the one before the failure.
In terms of reasons for the observed Kafka Streams’ fault recovery instabilities, the current understanding is that it is mostly due to which worker instance the failure is injected (e.g., if the consumer group leader is affected or not) and how the subsequent(s) repartitioning(s) are enacted. On the one hand, this outcome is aligned with Fault recovery observation 2. On the other hand, we believe that this potentially problematic behavior of Kafka Streams should be further characterized and improved.
Figure 6(c) shows Spark Structured Streaming’s behavior under recurrent failures. On the one hand, there are some minor noticeable fluctuations in the lag accumulated, latencies after failures (lowest latencies in the last failures injected), and throughput recovery times. These minor fluctuations are justified by the dynamic micro-batching and the fine-grain tasks mapping to the worker executor instances. On the other hand, the impact on performance and QoS is similar across the recurrent failures, which indicates that Spark Structured Streaming can also provide stable fault recovery.
Fault recovery observation 5 is that Flink and Spark Structured Streaming can provide stable fault recovery. Kafka Streams, on the contrary, has unstable behavior with volatile recovery time.
Figure 8 complements the results from Fig. 5. If we compare the recovery from the first failure shown in Fig. 5 to the five failures covered in Fig. 8, one can note that Flink and Spark Structured streaming recoveries are similar in all failures. Kafka Streams, on the contrary, result in failures having (once again) different impacts on its performance, mostly in terms of latency, lag accumulation, and lag reduction. We can then conclude that Fault recovery observation 2 also applies to worker instances crashes.
(a) Flink
(b) Kafka Streams
(c) Spark Structured Streaming
Figure 8. ALOS recovery: Two random pods executing worker instances crash for 5 minutes, repeated 5 times every 10 minutes.
8. Fault Recovery Time
The precise fault recovery time is another relevant metric that complements the results from Section 6 and Section 7. Considering that there is a lack of approaches and tools to automate the measurements of fault recovery times across many failures, we designed a simple detector based on moving average and standard deviation to automatically detect the failure as well as the time to recovery. The detector extracts the expected average and standard deviation for both throughput and latency based on the regular processing behavior that follows the warm-up period. The time frame ranged between 2 and 5 minutes, and 3 and 6 minutes, for throughput and latency, respectively. Figure 9 exemplifies the trace of automatically detected failures (of 2-pods kill) and their recoveries.
In short, if the moving average and standard deviation exceeded a given threshold (equal for all frameworks), then the detector would mark it as the point of failure injection. Afterward, another point of failure was strictly marked every 5 minutes. Similarly, the recovery time was detected when the estimators returned within the same threshold for a larger time frame, since once stability is achieved, throughput and latency should remain constant. Considering that the remaining experiments have a high correlation, for conciseness they are not shown here. We will provide the additional results and the parameters used for the thresholds in supplemental material.
Figure 9. Characterization of the automatic fault recovery times detection in the ALOS scenario of 2-pods kill.
Figure 10 summarizes the results of the fault recovery times. Noteworthy, throughput and latency (p90) recovery time are highly correlated. Flink achieves the fastest overall recovery time for both metrics. An exception is latency on 1-pod kill latency where Spark has a slightly lower median recovery ($m_{flink}=69.5$ seconds, $m_{spark}=68.5$ seconds) but at the cost of higher variability due to the observed spark’s fluctuations in throughput and latency.
Kafka Streams’ recovery times are longer and with higher variability, consistently with the discussion presented in Section 6. There are cases where full recovery was not achieved in the 5 minutes before the next failure was triggered, which is why the median can be high in some Kafka Streams’ results. Fault recovery observation 6 is that Flink and Spark Structured Streaming provide faster fault recoveries, and Flink demonstrates the highest level of stability.
Figure 10. Recovery times.
9. Concluding Remarks
Fault tolerance of data-intensive applications is of paramount importance in industry deployments. In this paper, we contribute with a more comprehensive evaluation method and fault recovery analysis. Such an evaluation was enabled by ShuffleBench’s (Henning et al., 2024) flexible deployment, parameters, and extension to fault recovery metrics.
9.1. Key Observations and Takeaways
We concluded that much has changed compared to the literature on how stream processing frameworks recover from failures. This motivates the need for benchmarking studies. The following are the key insights from the extensive experimental analysis:
  • Flink, Kafka Streams, and Spark Structured Streaming are resilient to different types and degrees of failures. However, the frameworks have different impacts on the applications running.
  • In use cases where at-least-once semantics guarantees are required, Flink provides currently the best recovery, which is both fast and stable. Spark Structured Streaming can also provide consistent fault recovery for applications where latency is not a critical requirement. Kafka Streams takes longer to recover from failures and presents volatile behavior.
  • Flink is the best candidate when exactly-once semantics guarantees are necessary. On the other hand, Kafka Streams has limited performance, and Spark Structured Streaming has limited native support for EOS.
9.2. Results Implications
The fault recovery capabilities should be a criterion for choosing a stream processing framework. Therefore, one of the main implications of our findings is that Kafka Streams, despite its growing interest, needs improvements in terms of fault recovery. This implication is relevant because our findings indicate that in contrast to what was reported in the past (van Dongen and van den Poel, 2021), Kafka Streams’ unstable recovery can compromise QoS and service level agreements. The finding that Kafka Streams faces performance instabilities after failures due to its current repartitioning strategy has relevant implications. We intend to share our findings with Kafka Streams’ community and contribute with future enhancements.
The positive recovery results seen in Flink imply that it is currently the optimal solution for what mostly motivates fault tolerance: resilience and continuous availability of services. Another relevant implication from the fault recovery stability measurements is that the common assumption that frameworks’ recovery performance is comparable before and after recovery (van Dongen and van den Poel, 2021) does not hold. In fact, our results indicate significant contrasts in recoveries across failures (see Section 8), which can impact the applications’ performance, QoS, and user experience.
With the significant impact of fault recovery on applications’ performance and QoS, we argue that modern performance benchmarking should be extended to support more realistic scenarios. We believe that performance is still a major goal, but it should be continuously improved. This would be a starting point to better representing real-world scalable computing deployments in clouds.
Fault recovery has a high impact on advanced optimizations necessary in modern stream processing, such as run-time self-adaptations for auto-scaling. This is because applying the runtime modifications involves recovery. Therefore, our findings are also directly applicable to auto-scaling. We also believe that our experimental method is generic enough to be useful for evaluating other domains. This is mostly because our method is comprehensive and includes insights into how chaos engineering can be applied to long-running computing systems.
9.3. Limitations and Threats to Validity
We expect all threats to the internal validity discussed in our previous works to also apply to our results (Henning et al., 2024). In addition to known threats to external validity, the fault measurements have additional configurations and executions threats. The failure injection interval was defined according to insights from real-world industry deployments, where there can be hundreds or thousands of worker instances, thus increasing failure likelihood. However, we still need more comprehensive statistical studies to provide insights into failure rates in large-scale cloud deployments. From our standpoint, we believe that new comprehensive studies are needed to characterize the failure rates in modern public hyperscalers, similar to what we have seen in high-performance computing centers (Schroeder and Gibson, 2009).
9.4. Future Work
We believe that future analyses to extend ours are still needed. A relevant lesson learned from the experiments is that 5 minutes between two injected failures is not enough to measure the recovery time in all scenarios. Therefore, future work is to experiment with failures injected over longer intervals. We also believe the following are necessary complementary analyses:
  • Cover other types of failures, such as in the manager node.
  • Extend our method for workloads with larger state sizes.
  • Evaluate the impact of the frameworks on the Kafka brokers’ performance, resource utilization, and costs.
  • Extended our analysis of fault recovery for more scenarios where EOS guarantees are needed.
  • Further analysis of the frameworks’ output correctness under failures with different fault tolerance semantic guarantees.
Acknowledgements.
We would like to thank Johannes Kepler University Linz and Dynatrace for co-funding this research. We thank Bernhard Kepplinger for providing insights on the method for failure detection.

### v2

@@
 Abstract.
-Nowadays, several software systems rely on stream processing architectures to deliver scalable performance and handle large volumes of data in near real time.
+Nowadays, several software systems rely on stream processing architectures to deliver scalable performance and handle large volumes of data in near real-time.
 Stream processing frameworks facilitate scalable computing by distributing the application’s execution across multiple machines.
-Despite performance being extensively studied, the measurement of fault tolerance—a key and most appealing feature offered by stream processing frameworks—has still not been measured properly with updated and comprehensive testbeds.
+Despite performance being extensively studied, the measurement of fault tolerance—a key feature offered by stream processing frameworks—has still not been measured properly with updated and comprehensive testbeds.
 Moreover, the impact that fault recovery can have on performance is mostly ignored.
@@
 Generally, our results indicate that much has changed compared to previous studies on fault recovery in distributed stream processing.
-In particular, the results indicate that Flink can be the fastest and stablest under failures.
-Moreover, Kafka Streams shows performance instabilities after failures, which is due to its current repartitioning strategy that can be suboptimal in terms of load balancing.
+In particular, the results indicate that Flink is the most stable and has one of the best fault recovery.
+Moreover, Kafka Streams shows performance instabilities after failures, which is due to its current rebalancing strategy that can be suboptimal in terms of load balancing.
 Spark Structured Streaming shows suitable fault recovery performance and stability, but with higher event latency.
@@
 benchmarking, fault tolerance, performance, stream processing
-copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software performanceccs: Computer systems organization Cloud computingccs: Information systems Stream management
 1. Introduction
 Stream processing refers to a computing paradigm and architecture pattern for building distributed, event-based software systems that process massive amounts of data in near real time (Hirzel et al., 2014).
-Today’s stream processing paradigms enable software engineers to model their applications’ business logic as high-level representations in a directed acyclic graph (DAG), without explicitly defining the physical execution plan.
+Today’s stream processing paradigms enable software engineers to model their applications’ business logic as high-level representations in directed acyclic graphs (DAGs), without explicitly defining the physical execution plan.
 This high level of abstraction is provided by industry-grade, open-source stream processing frameworks such as Spark (Armbrust et al., 2018), Flink (Carbone et al., 2015), or Kafka Streams (Wang et al., 2021).
@@
 They provide high-level APIs and domain-specific languages to define the processing logic as directed acyclic processing graphs that filter, transform, aggregate, and merge data streams.
-Over the last decade, several works have been published that evaluate the performance of distributed stream processing frameworks (Vogel et al., 2023; Henning and Hasselbring, 2024; Henning et al., 2024) or optimizations (Hirzel et al., 2014; Herodotou et al., 2020; Vogel et al., 2022).
-Beyond measuring performance metrics such as throughput and latency, fault tolerance stands as a critical requirement for continuously-operating production systems.
+Over the last decade, several works have been published that evaluate the performance of stream processing frameworks (Vogel et al., 2023; Henning et al., 2024) or optimizations (Hirzel et al., 2014; Herodotou et al., 2020).
+Beyond measuring performance metrics such as throughput and latency, fault tolerance is a critical requirement for continuously operating production systems.
 Failures can occur unpredictably across various levels, from physical infrastructure to software layers.
@@
 Particularly, this paper provides the following contributions to industry practitioners and the research community:
+  • An extension of ShuffleBench (Henning et al., 2024) for fault tolerance measurements in the different state-of-the-art stream processing frameworks, tooling to automate benchmarks’ execution, and data collection in Kubernetes-based cloud environments.
   • A comprehensive methodology for measuring fault recovery.
-The methodology covers stream processing’s most relevant metrics (Vogel et al., 2023), new metrics to improve fault recovery evaluations, a method for failure injection, and a method for automating fault recovery time measurements.
-  • An extension of ShuffleBench (Henning et al., 2024) for fault tolerance measurements in the different state-of-the-art stream processing frameworks, tooling to automate benchmarks’ execution, and data collection in Kubernetes-based cloud environments.
+The methodology covers stream processing’s most relevant metrics (throughput and latency) (Vogel et al., 2023) and new metrics to improve fault recovery measurements (fault recovery impact on performance, stability, and recovery time).
+  • A method inspired by chaos engineering that enables the injection of different types of recurrent failures in stream processing experiments.
+  • A method to automate fault recovery time measurements, available as supplementary material (Vogel et al., 2024b).
   • An experimental analysis covering application metrics, system resources utilization, and fault recovery.
-[^1] On publication of this study, we will also provide as supplemental material an extended replication package of ShuffleBench for fault tolerance measurements.
-Moreover, the experimental data is also to be shared such that other researchers and practitioners may repeat and extend our work.
 Our evaluation covers the open-source stream processing frameworks Flink, Kafka Streams, and Spark Structured Streaming due to their industry acceptance and academic relevance (Vogel et al., 2023).
+The experimental data and an extended replication package of ShuffleBench (Henning et al., 2024) for fault tolerance measurements are available as supplementary material (Vogel et al., 2024b).
 2. Background
@@
 2.1. Distributed Stream Processing
-Stream processing frameworks perform operations such as filterings, transformations, or aggregations in near real-time on continuous streams of data (Hirzel et al., 2014).
-State-of-the-art frameworks are designed for high throughput and low-latency processing, while also scaling with massive amounts of data (Fragkoulis et al., 2024; Henning and Hasselbring, 2024).
-To address these requirements, they run in a distributed fashion on commodity hardware.
+Stream processing frameworks perform operations such as filterings, transformations, or aggregations in near real time on continuous streams of data (Hirzel et al., 2014).
+State-of-the-art frameworks are designed for high throughput and low-latency processing, while also scaling with massive amounts of data (Fragkoulis et al., 2024; Henning et al., 2024).
+To address these requirements, they run in a distributed fashion on commodity hardware, nowadays often in managed cloud environments.
 A key advantage of stream processing frameworks is that they provide dataflow models that abstract aspects such as cluster management, state management, and time semantics from their users (Sax et al., 2018).
-With such models, engineers describe the processing logic in DAGs of processing operators.
 The frameworks allow the initiation of multiple worker instances across various compute nodes with multiple threads, each instance handling a distinct portion of the data.
@@
 In the absence of fault tolerance, stream processing applications would be forced to restart data processing from scratch whenever there is a loss due to a failure.
-Given that applications may have varying fault tolerance needs, the stream processing domain focuses on different processing semantics that precisely describe how a system is affected by failures (Fragkoulis et al., 2024).
-The relevant terms to characterize the fault tolerance semantics guarantees are at-most-once semantics (AMOS), at-least-once semantics (ALOS), and exactly-once semantics (EOS).
-Currently, the majority of stream processing applications demand reliable executions with strong semantics guarantees.
-Thus, AMOS is obsolete (Fragkoulis et al., 2024) and we focus on ALOS and EOS.
+Processing semantics guarantees describe how a system is affected by failures.
+Various guarantees are provided to accommodate the varying fault tolerance needs of different use cases (Fragkoulis et al., 2024).
+Relevant terms are at-most-once semantics (AMOS), at-least-once semantics (ALOS), and exactly-once semantics (EOS).
+Given that the majority of stream processing applications require reliable executions with strong semantics guarantees, we have shifted our focus from AMOS, which can be considered obsolete, to ALOS and EOS (Fragkoulis et al., 2024).
 At-least-once semantics
@@
 Therefore, EOS requires additional implementation and has a performance impact on records processing (van Dongen and van den Poel, 2021).
-In this paper, we focus on semantics guarantees regarding the system output because this is the most relevant for the majority of use cases and also covers the guarantees in the internal state updates.
+In this paper, we focus on the semantic guarantees of the system output, as these are most relevant for the majority of use cases and also encompass the guarantees for internal state updates.
 The stream processing frameworks provide different recovery semantic guarantees.
 This is important because the required guarantee varies according to the stream processing application and its use cases.
-In some scenarios, guaranteeing exactly-once can be of paramount importance, e.g., avoiding duplicate alerts in an observability platform.
-In other cases, at-least-once with rarely duplicated outputs is not equally detrimental, e.g.
-a monitoring system that receives twice the same CPU metric of a timestamp.
+In some scenarios, guaranteeing exactly-once delivery is of paramount importance, such as avoiding duplicate alerts in an observability platform.
+In other cases, at-least-once with rarely duplicated outputs is not equally detrimental, e.g., an information dashboard that receives twice the same CPU metric of a timestamp.
 We intend to measure the impacts of the stream processing frameworks when running with the fault recovery guarantees.
@@
 3.1. State-of-the-art
-Recent publications have presented various experimental evaluations of fault recovery in stream processing (Vogel et al., 2023).
+Recent publications provide experimental evaluations of fault recovery in stream processing (Vogel et al., 2023).
 Although many studies consider failures of worker instances (Lu et al., 2014; Qian et al., 2016; Lopez et al., 2016), the most updated and relevant is (van Dongen and van den Poel, 2021).
-It dives into the implementation, performance, and efficiency of fault recovery in four state-of-the-art stream processing frameworks: Spark Streaming, Flink, Spark Structured Streaming, and Kafka Streams.
-In addition, they evaluated the behavior of these frameworks under different types of faults and settings, including master failure with and without high-availability setups, driver failures for Spark frameworks, worker failure with or without exactly-once semantics, and application failures.
+It dives into the implementation, performance, and efficiency of fault recovery in four stream processing frameworks: Spark Streaming, Flink, Spark Structured Streaming, and Kafka Streams.
+In addition,  van Dongen and van den Poel (2021) evaluate the behavior of these frameworks under different types of faults and settings, including master failure with and without high-availability setups, driver failures for Spark frameworks, worker failure with or without exactly-once semantics, and application failures.
 Wang et al.
@@
 Wang et al.
-(2021) explain Kafka Streams’ processing logic and demonstrate how Kafka Streams behaves in practice with large-scale deployments and performance insights exhibiting its flexible and low-overhead trade-offs.
-A very recent paper from Siachamis et al.
-(2024) focuses on complementary aspects of fault tolerance.
-The authors explored three approaches for checkpoint protocols (coordinated, uncoordinated, and communication-induced) aimed at fault tolerance in stream processing.
-The analysis delved into the theoretical strengths and weaknesses inherent in each approach.
-Their findings not only validated the prevailing preference for the coordinated approach but also illustrated instances, where the uncoordinated approach exhibited resilience and better performance.
-Considering that many optimizations can be done for fault tolerance (Hirzel et al., 2014; Vogel et al., 2022), Su and Zhou (2021) address the issue of recovering from correlated failures in large-scale clusters.
+(2021) demonstrate how Kafka Streams works on large-scale deployments and performance insights exhibiting its flexible and low-overhead trade-offs.
+A recent paper from Siachamis et al.
+(2024) focuses on complementary aspects of fault tolerance, exploring three approaches for checkpoint protocols (coordinated, uncoordinated, and communication-induced) aimed at fault tolerance in stream processing.
+The analysis delves into the approaches’ theoretical strengths and weaknesses.
+Their findings validate the preference for the coordinated approach and illustrated instances where the uncoordinated approach provides performance gains.
+Considering that many optimizations can be done for fault tolerance (Hirzel et al., 2014), Su and Zhou (2021) address the issue of recovering from correlated failures in large-scale clusters.
 The authors propose an incremental and query-centric recovery paradigm.
@@
 3.2. Research Gaps
-While fault recovery in stream processing has been addressed in various related studies, our understanding is that the methods and insights provided offer limited coverage.
-Despite being recognized as essential within a leading cloud observability platform, we identified significant research and analysis gaps in the state-of-the-art regarding fault recovery (Vogel et al., 2023):
-  • Considering the continuous development of fault tolerance features and optimizations by the frameworks’ open-source communities, there is a need for updated measurements on frameworks’ executions facing worker instances’ failures.
-  • The few studies that compare the frameworks regarding worker instances’ failures cover a single failure (van Dongen and van den Poel, 2021; Wang et al., 2022).
-However, a large physical machine crash impacts several smaller virtual machines or pods running within it.
+While fault recovery in stream processing has been addressed in various related studies, our understanding is that the methods and insights provided in the literature offer limited coverage.
+We believe that the following are relevant research gaps in the state-of-the-art regarding fault recovery on distributed stream processing:
+  1. (1)
+    Considering the continuous development of fault tolerance features and optimizations by the frameworks’ open-source communities (Fragkoulis et al., 2024), there is a need for updated measurements on frameworks’ executions facing worker instances’ failures.
+  2. (2)
+    Few studies compare how different frameworks recover worker instances and those that do focus on failures affecting a single instance (van Dongen and van den Poel, 2021; Wang et al., 2022).
+However, a large physical machine crash impacts several virtual machines or pods.
 What can happen is that more than one worker instance fails.
-However, it remains unclear what is the impact of such a more critical failure in the running stream processing applications.
-  • The existing literature only considers that worker instances can fail a single time (van Dongen and van den Poel, 2021; Wang et al., 2022).
-In real-world industry deployments, failures happen recurrently and at unpredictable times.
-One can assume that the frameworks’ recovery will always be the same, but we observed that such an assumption can have weak coverage with the reality of large deployments.
+However, the impact of such a more critical failure on running stream processing applications remains unclear.
+  3. (3)
+    The existing literature only covers scenarios where failures occur a single time (van Dongen and van den Poel, 2021; Wang et al., 2022).
+However, in real-world industry deployments, failures happen recurrently and unpredictably.
+One can assume that the frameworks’ recovery will always be the same, but we observed that such an assumption could have weak coverage with the reality of large deployments.
 This occurs mostly because the impact that a given failure has in the applications varies according to the exact moment that the failure happens (e.g., how far from the last saved state checkpoint), what working instances are affected by the failure, and the effectiveness of the subsequent recovery (Geldenhuys et al., 2022).
-Therefore, the level of fault recovery stability across recurrent failures remains unclear.
+Therefore, the level of fault recovery stability across recurrent failures demands more research.
 4. Experimental Setup
@@
 It comes as a ready-to-use open-source software [^1] https://github.com/dynatrace-research/ShuffleBench utilizing existing Kubernetes tooling and providing implementations for different state-of-the-art stream processing frameworks.
-ShuffleBench utilizes and extends the benchmarking framework Theodolite (Henning and Hasselbring, 2022) to automate the benchmark execution in Kubernetes-based cloud environments.
+ShuffleBench adopts and extends the benchmarking framework Theodolite (Henning and Hasselbring, 2022) to automate the benchmark execution in Kubernetes-based cloud environments.
 Figure 1 depicts our dataflow architecture for a corresponding stream processing application.
-In short, data records are read from a messaging system (Kafka in our case), assigned to keys, and shuffled such that all records having the same key are forwarded to the same worker instance.
-There, a key-specific custom aggregation is performed and, optionally, an output event is created that is then written back to Kafka.
-We refer to our previous work for further details about ShuffleBench (Henning et al., 2024).
+Data records are read from a messaging system (e.g., Kafka), assigned to keys (zero to many), and shuffled such that all records having the same key are forwarded to the same worker instance.
+There, a key-specific custom aggregation is performed, and an output event is created that is then written back to Kafka.
+We refer to our previous work for a detailed description of ShuffleBench (Henning et al., 2024).
 Figure 1. The ShuffleBench dataflow architecture of a stream processing application with three worker instances (Henning et al., 2024).
@@
 In this work, we extend ShuffleBench for fault tolerance measurements.
-Figure 2 provides an overview of ShuffleBench’s benchmark components and their extension to fault tolerance, where the scenarios and metrics are described in Section 5.
-We conduct our experimental evaluation in a Kubernetes cluster managed by the Elastic Kubernetes Service of Amazon cloud with the following pods running in cloud instances: 2 load generators (m6i.xlarge instances), 2 Kafka brokers (m6i.2xlarge), a manager node used on Flink and Spark (m6i.large), and 8 pods for worker instances with 1 CPU, 3 GB of RAM (running on c5.large instances).
-In our experiments, we set up 20 000 real-time consumers that all have the same selectivity, which sum up to 50 %, meaning that each record is forwarded to $0.5$ consumers on average.
+Figure 2 provides an overview of ShuffleBench’s components and their extension for fault tolerance (see the scenarios and metrics in Section 5).
+We conduct our experimental evaluation in a Kubernetes cluster managed by the Elastic Kubernetes Service of Amazon Web Services with the following pods running in dedicated cloud instances: 2 load generators (m6i.xlarge instances), 2 Kafka brokers (m6i.2xlarge), a manager node used on Flink and Spark (m6i.large), and 8 worker instances pods with 1 CPU, 3 GB of RAM (running on c5.large instances).
+In our experiments, we set up 20 000 real-time consumers that all have the same selectivity, which sum up to 50 %, meaning that each record is forwarded to $0.50.5$ consumers on average.
 Each consumer emits an output event for every record received.
@@
 Fault tolerance on Flink and Spark Structured Streaming involves periodic state backup via checkpointing, demanding a reliable filesystem to store state copies (van Dongen and van den Poel, 2021).
-We used in our setup an Amazon Elastic File System (EFS) filesystem volume mounted in the pods.
+We used in our setup an Amazon Elastic File System (EFS) volume mounted in the pods.
 Kafka Streams provides fault recovery with its strong integration with Kafka.
@@
 5.2. Failure Injection Scenarios
-Experimental scenarios were inspired by chaos engineering (Basiri et al., 2016).
-Failures are injected using Chaos Mesh, [^3] https://github.com/chaos-mesh/chaos-mesh an open-source chaos engineering platform integrated with the Kubernetes deployment.
-We consider two scenarios:
+Our experimental scenarios were inspired by chaos engineering (Basiri et al., 2016) using Chaos Mesh [^2] Chaos Mesh is an open-source platform: https://github.com/chaos-mesh/chaos-mesh.
+to inject failures in two scenarios:
 Pods kill:
 Random pods executing worker instances are deleted.
-When the pods are no longer accessible, both the records they were processing and their associated state are lost.
-Therefore, the frameworks have to trigger recovery from such a failure.
+When the pods are no longer accessible, both the records they were processing and their internal state are lost.
+Therefore, the frameworks trigger recovery from such a failure.
 Deployed with Kubernetes, new pods will be created and added to the pool of worker instances within an interval of seconds (usually taking from 2 to 10 seconds).
@@
 Such values were determined to simulate simple and critical failure scenarios.
-Such failures are recurrently injected every 5 minutes to evaluate the fault recovery stability.
+The failures are recurrently injected every 5 minutes to evaluate the fault recovery stability.
 This interval was defined for a reasonable recovery time between failures and also to avoid unnecessary long executions that would waste resources.
@@
 Then, executions can resume the normal state.
-Moreover, 5 minutes is a suitable value for creating and adding new virtual machines (Su and Zhou, 2021) to the pool.
+Moreover, 5 minutes is a suitable value for creating and adding new virtual machines to the pool (Su and Zhou, 2021).
 5.3. Metrics
@@
 A higher throughput is usually desired in many use cases.
-We measure input throughput refers to consuming from the Kafka input topic, and output throughput which is the application output written into another Kafka topic.
-We also measure event latency (the lower latency tends to be better) as an average of a 10-second window.
-We show latency percentiles to measure the distribution of events times: the p50 (50th percentile), p90, and p99.
+We measured input throughput (the rate of records consumed from the Kafka input topic, see Fig.
+1) and output throughput (the rate of records written to the Kafka output topic).
+We also measure event latency (lower is better) as an average of a 10-second window.
+We show latency percentiles to measure the distribution of events times: the median (p50), 90th percentile (p90), and 99th percentile (p99).
 We also measure the consumer lag corresponding to the number of records buffered in the input topic.
 The fault recovery on stream processing is measured with the metrics: Fault recovery performance, the impact of fault tolerance guarantees, stability, and fault recovery time (see Figure 2).
-We also present the following system monitoring indicators collected every 2 seconds from the cloud instances that run the worker pods: CPU utilization, memory utilization, and network traffic.
-[^4] To enhance visual clarity, the CPU utilization and network traffic of worker nodes are displayed using a 10-value moving window average.
+We also present the following system monitoring indicators collected every 2 seconds from the cloud instances that run the worker pods: CPU usage, memory utilization, and network traffic.
+[^3] To enhance visual clarity, the CPU usage and network traffic of worker nodes are displayed using a 10-value moving window average.
+The method described here is applied in Sections 6, 7 and 8 to conduct measurements and extract insights.
+In this paper, we show the most insightful results, complementary results can be found in the supplementary material (Vogel et al., 2024b).
 6. Fault Recovery Characterization
-This section shows results demonstrating how the injected failures impact the frameworks’ execution as well as how they react to the failures.
-We provide results on the application metrics and system resource utilization (see more in Section 5.3).
-This section intends to demonstrate the potential impacts of faults and their recovery on the different stream processing frameworks and their Quality of Service (QoS).
-In this section, we show first failures for the sake of visual clarity, Section 7 extends these characterizations providing results with up to 10 recurrent failures injected.
+This section demonstrates how the injected failures impact the frameworks’ executions.
+The results cover application metrics and resource utilization (see more in Section 5.3).
+This section intends to demonstrate the potential impact of faults, their recovery on the different stream processing frameworks, and their Quality of Service (QoS).
+For visual clarity, The figures and plots provided for visual inspection depict either (i) the first three failures injected in the case of pod kills and (ii) the first failure followed by the frameworks’ response in the case of pod crashes.
+Section 7 extends the characterization with results from up to 10 injected failures.
 6.1. ALOS Recovery Characterization
@@
 Figure 3. ALOS fault recovery characterization: Random pods executing worker instances are killed, repeated 3 times.
-Figure 3 shows how the executions with the different frameworks behave when pods executing worker instances are killed (represented by the red line).
-To have more comprehensive characterizations, we present results with three failures that occur periodically with an interval of 5 minutes.
+Figure 3 shows executions with the different frameworks with pods executing worker instances being killed (represented by the red line).
 It is also important to note that the plots show different metrics (see y-axis labels) over time.
-When the failures occur, the worker instance impacted is lost.
-Then, the Kubernetes infrastructure creates a new pod that is added to the pool of worker instances.
-Moreover, the stream processing frameworks are expected to start scheduling computations to new worker instances.
-This is achieved in Flink and Kafka Streams by rebalancing the partitions to worker instances.
-In Spark Structured Streaming, the worker instances are added to the pool of executors and they start receiving tasks to compute.
+When the failures occur, the worker instance affected is lost.
+Then, the Kubernetes creates a new pod, which is added to the pool of worker instances.
+Moreover, the stream processing frameworks start scheduling computations to the new worker instances.
+This is achieved in Flink and Kafka Streams by rebalancing the partitions among the worker instances.
+In Spark Structured Streaming, the new worker instances start receiving tasks to compute.
 Fault recovery observation 1: The frameworks recover correctly from the failures.
-From a quality of services perspective, it is noticeable that Flink executions recover quicker from failures in comparison to Kafka Streams and Spark Structured Streaming, which is evinced by a shorter downtime in the output throughput and quicker reduction of the latencies.
-Performance observation 1: After failures, Flink can present fluctuations in the data consumed from Kafka.
-To the best of our knowledge, this phenomenon has not been previously reported.
-The variation in the amount of data ingested after failures may be attributed to the uncertainty of the timing of failures relative to the previously saved checkpoints.
-However, we have deferred further exploration of this phenomenon to future work, as it does not significantly impact the output throughput or the latency.
-Performance observation 2: Kafka Streams achieves a stable throughput with and without failures of one worker instance pod, but after failure has fluctuations in resources usage.
-Kafka Streams is the one that accumulates more lag after failures, which is mostly due to its default partition assignment strategy triggered after the failures.
-After the first failure, there are fluctuations in the CPU utilization and network traffic in the worker machines, which indicates a potential imbalance.
-This also relates to the partition assignment strategy where some worker instances in the consumer group are assigned to more partitions or partitions with more lag accumulated.
-See more in the resource consumption observation 2.
-Performance observation 3: Spark Structured Streaming’s dynamic definition of micro-batch intervals is finding suitable configurations.
-Dynamic micro-batching is evident in the first few minutes of the execution with the throughput and lag fluctuating significantly.
-Such fluctuations happen because the micro-batch interval is too large causing more data to be consumed and batches being triggered less frequently.
-After a few minutes, the micro-batch interval is self-regulated to close to 2 seconds, which is a configuration suitable to sustain the input throughput and maintain a reasonable latency (the p50 latency is comparable to the micro-batch interval).
+However, when comparing the frameworks, it is noticeable that Flink executions recover quicker from failures.
+This observation is evinced by a shorter downtime in the output throughput and a faster latency recovery.
+Performance observation 1 is that after failures, Flink occasionally shows fluctuations in the input throughput.
+To the best of our knowledge, this observation has not been previously reported.
+Such fluctuations in the input throughput after failures can be due to the variation of the failures instant relative to the last saved checkpoints.
+Considering that these variations do not significantly impact the output throughput or the latency, further analysis is left for future work.
+Performance observation 2: Kafka Streams achieves a stable throughput with and without failures of one worker instance pod, but Kafka Streams’ executions are subject to unstable latencies.
+The Kafka Streams’ partition assignment strategy triggers a rebalance after failures, causing its executions to accumulate more lag.
+Rebalancing significantly increases event latency and shows fluctuations in the worker machines’ CPU usage and network traffic, indicating a potential imbalance.
+In fact, Kafka Streams’ performance following failures is mostly impacted by its partition assignment strategy, which unevenly distributes partition to the consumer group’s worker instances.
+Some worker instances may be assigned to more partitions or partitions with more accumulated lag.
+Then, some workers become overloaded while others are underutilized.
+[^4] Further analysis and potential optimizations in our follow-up paper (Vogel et al., 2024a).
+Performance observation 3: Spark Structured Streaming’s dynamic micro-batch intervals stabilize after a few minutes on suitable configurations.
+Dynamic micro-batching is noticeable in the first few minutes of execution due to fluctuations in throughput and lag, caused by an overly large micro-batch interval that results in more data being consumed and less frequent batches being triggered.
+Spark’s self-regulated micro-batch interval stabilizes at around 2 seconds per batch after a few minutes of execution.
+This configuration sustains the input throughput while maintaining reasonable latency.
+In our experience, the measured p50 latency is comparable to the micro-batch interval.
 Still, Spark’s latency is significantly higher than the other frameworks that process the records based on events instead of micro-batches.
-This outcome is consistent with the related literature (van Dongen and van den Poel, 2021).
+This insight is consistent with the related literature (van Dongen and van den Poel, 2021).
 (a) Fault recovery
@@
 Figure 4. ALOS recovery characterization: Two random pods executing worker instances are killed, repeated 3 times.
-Resource consumption observation 1: Spark Structured Streaming has the most dynamic CPUs utilization.
-This is due to its modern execution model that processes based on fine-grain tasks assigned to the spark’s executors (worker instances).
-Tasks are created within the so-called stages that execute the micro-batches, the number of such tasks is determined based on the number of partitions (consumed from Kafka and subsequent shuffle partitions between DAG’s stages).
-Such a dynamic execution model can be more scalable and suitable for effective load-balancing but at the cost of potential overheads of having to assign all the fine-grain tasks to the executors.
-It is also noticeable that, from the operating system view, Spark Structured Streaming execution is the one that allocates the highest memory space.
-[^5] This memory observation should not be taken as evidence because the way the Java virtual machine allocates memory and computes can vary across the frameworks.
-Therefore, the operating system view of the memory usage does not guarantee accuracy.
-Considering that memory consumption can have a heavy impact on performance and cost, this motivates future analysis to comprehend the root causes and potential mitigations.
-One direction is to measure the garbage collection (GC) effectiveness and performance, where the Spark’s GC was fine-tuned in related works (van Dongen and van den Poel, 2021).
+Resource consumption observation 1: Spark Structured Streaming has the most dynamic CPUs usage.
+This is due to its modern execution model based on fine-grain tasks assigned to Spark’s executors (worker instances).
+Tasks are created within the so-called stages that execute the micro-batches.
+The number of tasks is defined based on the number of partitions (consumed from Kafka and shuffle partitions between the DAG’s stages).
+Such a dynamic execution model can be more scalable and provide load balancing, but it comes at the cost of potential scheduling overheads due to assigning all the fine-grained tasks to executors.
+From the operating system view, Spark Structured Streaming execution is the one that allocates the largest memory space.
+[^5] Considering that how Java Virtual Machine allocates memory and computes can vary across the frameworks, this memory observation should not be taken as evidence because the operating system’s view of memory usage does not guarantee accuracy.
+Considering that memory consumption can heavily impact performance, this motivates future analysis to comprehend the root causes and potential mitigations.
+One direction is to measure the garbage collection (GC) effectiveness and performance, where Spark’s GC was fine-tuned in related work (van Dongen and van den Poel, 2021).
 Moreover, this behavior could be a potential reason for Spark’s performance limits found in our previous work (Henning et al., 2024).
 Resource consumption observation 2: From a resource consumption perspective, Kafka Streams shows unstable behavior after the failures are injected.
-This is supported mostly by varied CPU utilization and network traffic in the worker nodes, which suggests some imbalance that we believe to be load-related when repartitioning after the failures.
-This outcome aligns with performance observation 2: the partition assignment strategy can be suboptimal after failures making some workers overloaded and others underutilized.
-Furthermore, from an operating system view, Kafka Streams execution is the one allocating least memory, which is a positive outcome from a resource efficiency view.
+This is supported mostly by varied CPU usage and network traffic in the worker nodes, which suggests some imbalances aligned with performance observation 2.
+Furthermore, Kafka Streams’ executions allocate the least memory, which is a positive outcome from a resource efficiency view.
 Resource consumption observation 3: Flink shows consistent resource utilization under failures.
-There are variations caused by the failures, but the execution returns to a normal stable state.
-It is worth noting that Flink’s execution is the one causing more CPU utilization in the Kafka Broker servers, which are already large-sized instances.
-This may be attributed to Flink’s fine-grained event-based processing, which continuously generates outputs to the Kafka topic.
-We noted that Flink’s performance may rely more than other frameworks on the behavior of the Kafka brokers and expect to address this in a future analysis.
-Figure 4 shows the scenario where periodically two random pods executing worker instances are killed.
-We expect this scenario to be more challenging as more workers failing means that more internal state can be lost and more records need to be replayed from the input source and reprocessed.
+Although the failures cause fluctuations, the executions return to a normal stable state.
+It is worth noting that Flink’s execution is the one causing more CPU usage in the large instances of Kafka Broker servers.
+This can be due to Flink’s fine-grained event-based processing, which continuously generates outputs to the Kafka topic.
+We noted that Flink’s performance may rely more than other frameworks on the behavior of the Kafka brokers.
+We intend to explore this insight further in a future analysis.
+Figure 4 shows the scenario where two random pods are periodically killed.
+We expect this scenario of more workers failing to be more challenging as more internal state is lost and more records need to be reprocessed by being replayed from the input source.
 Compared to Fig.
@@
 However, it is noticeable in Fig.
-4 that as Kafka Streams’ execution accumulates more lag after the failures, its event latency increases dramatically.
+4 that Kafka Streams’ execution accumulates more lag after the failures and its event latency increases dramatically.
 (a) Fault recovery
@@
 Figure 5. ALOS fault recovery characterization: Two random pods executing worker instances crash for 5 minutes.
-With a more critical failure as the one with two pods failing, the unbalances caused by Kafka Stream’s repartitions after failures evinced in Performance observation 2 and Resource consumption observation 2 become more critical and impactful.
-On the one hand, Kafka Streams strategy can evenly distribute the partitions to the available workers in some cases of repartitions after failures.
+Kafka Streams’ execution shows good performance after failures in some cases, e.g., after the third failure demonstrated in Fig.
+4 where the rebalance enables the lag to be processed quickly and the latency reduced.
+Figure 4(b) shows that a Kafka Streams’ worker node (indicated in green) consumes more CPUs and has higher network traffic while the lag is being reduced.
+This indicates that a worker instance pod on this specific node was assigned to a lagged partition and processed the lagged records.
 On the other hand, there are many cases where the distribution of partitions is suboptimal.
-For instance, the results from Fig.
-4 indicate that after the first failure, the repartition was suboptimal with some workers taking too long to process the lag accumulated.
-However, after the third failure, the repartition seems to be optimal in such a way that the lag is fast processed and the latency reduced.
-This is supported by the resource utilization data that shows that a worker node (green color) consumes more CPUs and has higher network traffic while the lag is being reduced.
-Potentially this specific node was running a pod of a worker instance assigned to a lagged partition that was able to process the lagged records.
-Therefore, Fault recovery observation 2 is that Kafka Streams’ fault recovery can have an inconsistent impact on the applications’ performance across recurrent failures.
-Although four worker instances failure is a more critical scenario, the results are aligned with the insights from the failures of two worker instances.
-Therefore, for the sake of conciseness, we do not show here results with 4 pods being killed.
-Figure 5 shows the scenario where the failure injected is that random pods executing worker instances face an operating system crash (see Section 5.2 for further details and motivations for this scenario).
-As they become unavailable, the records being processed by them and their local state is lost.
-In contrast with the pods killed failure, in this case, no new pods are created.
-Therefore, the execution has to continue processing with fewer worker instances.
-Focusing on Fig.
-4(a) after the failure is injected, in Fig.
-4(b) significant drops in system utilization are noticeable in the machines executing the two crashed pods.
-On the other hand, the active worker instances in all three frameworks increase their resource utilization to cope with the application processing demand.
-When the failure injected is halted (the gray line), the executions can return to the normal state with crashed worked instances being re-added to the workers pool.
+For example, Fig.
+4 shows that after the first failure, some workers take too long to process the accumulated lag compromising the recovery.
+Therefore, Fault recovery observation 2 is that Kafka Streams fault recovery’s impact on application performance varies across recurrent failures.
+We do not show here the results from four pods being killed because they are aligned with the insights from the failures of two worker instances (Vogel et al., 2024b).
+Figure 5 shows the scenario of failure injected where random pods executing worker instances crash (see more in Section 5.2).
+When pods become unavailable, the local state is lost.
+The execution must continue processing with fewer worker instances, as new pods are not created when pods crash.
+Significant drops in system utilization are noticeable after the failure in the machines executing the two crashed pods Fig.
+5(b).
+Moreover, in the three frameworks, the remaining active workers increased their resource utilization to cope with the failure.
+When the injected failure is halted (the gray line), the executions can return to the normal state with crashed worked instances being re-added to the worker pool.
 From the performance and QoS standpoint evinced in Fig.
-4(a), Fault recovery observation 3 is that the executions of the three frameworks recovered from a critical failure that reduces the resource availability, and when the failure is halted they returned to a comparable state as the one before the failure.
-One can note that Flink’s execution is the one that recovers faster to a normal performance state.
-Kafka Streams is the one accumulating more lag and its execution took almost 3 minutes to normalize the latency.
-Spark Structured Streaming demonstrates fluctuations after the failure, which is due to reprocessing of the tasks that were assigned to the failed executors and due to the dynamic micro-batching.
-With fewer executors and consequently fewer worker instances resources, Spark Structured Streaming stabilized the micro-batching interval with a higher value of around 3 seconds.
-The increase in the micro-batching interval was seemingly necessary to cope with the input load speed.
-However, this results in a slightly higher latency.
-Moreover, it is noticeable that Spark Structured Streaming’s output throughput faces more fluctuations after the failure is injected.
-After the failure is halted the executions with the three frameworks return to normal states.
-Comparing the different worker instances, Kafka Streams ends up with some noticeable imbalances in network traffic and CPU utilization (Fig.
-4(b)).
+5(a), Fault recovery observation 3 is that the executions of the three frameworks recovered from a critical failure with fewer resource available, they also returned to a comparable state as the one before the failure when the failure is halted.
+One can note in Fig.
+5(a) that Flink’s execution is the one that recovers faster.
+Kafka Streams is the one accumulating more lag and its execution took almost 3 minutes to reduce the latency.
+After the failure is halted, Kafka Streams ends up with some noticeable imbalances in network traffic and CPU usage (Fig.
+5(b)).
+Figure 5(a) shows Spark Structured Streaming with fluctuations after the failure, which is due to the reprocessing of the failed executors’ tasks and its dynamic micro-batching.
+The micro-batching interval was increased to a higher value of around 3 seconds, attempting to cope with the same data volume and fewer processing resources.
+However, this caused a slight increase in Spark Structured Streaming’s latency and fluctuations in the output throughput.
 6.2. EOS Recovery Characterization
 Figure 6. EOS recovery characterization: Random pods executing worker instances are killed, repeated 3 times every 5 minutes.
-Supporting exactly-once semantics (EOS) demands further configurations and deployments.
-One more Kafka broker was necessary for resource stability and the fault tolerance guarantee had to be configured in the frameworks.
-Moreover, as also done in the related literature (van Dongen and van den Poel, 2021; Wang et al., 2021) when testing EOS, the configuration of the isolation level is set to read only committed data, which prevents the reading of duplicated messages in downstream operators.
-Considering that Spark Structured Streaming can not natively guarantee exactly-once when writing the output to a Kafka topic, we adopt the same strategy of the related literature (van Dongen and van den Poel, 2021) and do not present Spark Structured Streaming with exactly-once.
-For conciseness, we show only the results with one pod kill that already provides the insights necessary to highlight Flink’s advantages and Kafka Streams’ shortcomings when EOS is needed.
+Supporting exactly-once semantics (EOS) requires additional deployment considerations.
+One more Kafka broker was needed for resource stability and the fault tolerance guarantee had to be enabled in the frameworks.
+Moreover, the isolation level was configured to read only committed data to prevent the reading of duplicated messages in downstream operators, consistent with practices in related literature (van Dongen and van den Poel, 2021; Wang et al., 2021).
+Considering that Spark Structured Streaming can not natively guarantee exactly-once when writing the output to a Kafka topic.
+This relates to the output commit problem solved in Flink using the two-phase commit protocol and in Kafka Streams with custom sinks for rollback-recovery (Carbone et al., 2017; Fragkoulis et al., 2024).
+We do not show results from Spark Structured Streaming on exactly-once experiments, consistent with practices in related literature (van Dongen and van den Poel, 2021).
+We show only the results with one pod kill that already provides the insights necessary to highlight Flink’s advantages and Kafka Streams’ shortcomings when EOS is needed.
 In Fig.
-6 it becomes evident that Flink, compared to Kafka Streams, provides better performance and fault recovery, [^6] We do not show the results from machine system monitoring of resources because it becomes repetitive and without new insights w.r.t.
-the previous results.
+6 it becomes evident that Flink provides better performance and fault recovery than Kafka Streams, [^6] System monitoring results are not included in this section for the sake of conciseness, as there are no new insights from a resources perspective.
 thus resulting in Fault recovery observation 4: Flink provides better performance and fault recovery for use cases that demand exactly-once guarantees.
-Focusing on the output throughput, both frameworks’ executions are similar until the first failure occurs.
-Then, Flink recovers faster in less than a minute to a normal output throughput that becomes stable.
-On the other hand, Kafka Streams took more than one minute to reach a necessarily higher output throughput.
-Moreover, after the failure, Kafka Streams’ output throughput is unstable and subject to fluctuations.
-We believe that this is due to the lag accumulated and being too slowly processed.
-This aligns with Performance observation 2 and Fault recovery observation 2.
-However, the scenario with exactly once is a more critical one, which results in a stronger impact of Kafka Streams’ suboptimal implementations.
-One can note that Flink achieved a much lower latency in Fig.
-6.
-Kafka Streams stabilized in a latency above 2 seconds and after the first failure, the latency does not normalize due to the large amount of lagged records.
-After a failure, Kafka Streams’ latency exceeds the 10-second metric bucket limit (Henning et al., 2024), rendering it invisible in the y-range plotted.
-These observations in Kafka Streams are, to the best of our knowledge, novel findings w.r.t.
-related approaches covering Kafka Streams with EOS (van Dongen and van den Poel, 2021; Wang et al., 2021).
-This can be mostly due to the fact the Kafka Streams was only tested under failures in (van Dongen and van den Poel, 2021) with a single failure injected.
-Moreover, technical reasons for the higher latency are due to the records buffering in Kafka Streams’ repartition topic and that its partition assignment strategy has been changed since (van Dongen and van den Poel, 2021), which implies that the current strategy is suboptimal for rebalancing after failures.
-7. Fault Recovery Stability
-This section presents the results of fault recovery stability under recurrent failures injected.
-Note the different ranges and scales on the y-axis.
-For conciseness, we limit the discussion to the results with ALOS guarantees as those are the most complete and insightful.
 (a) Flink
@@
 Figure 7. ALOS fault recovery: Random pods executing worker instances are killed, repeated 10 times every 5 minutes.
+(a) Flink
+(b) Kafka Streams
+(c) Spark Structured Streaming
+Figure 8. ALOS recovery: Two random pods executing worker instances crash for 5 minutes, repeated 5 times every 10 minutes.
+One can note that Flink achieved a much lower latency in Fig.
+6.
+Kafka Streams stabilized in a latency above 2 seconds and the latency does not normalize after the failure due to the significant lag.
+After failures, Kafka Streams’ latency is invisible in the y-range as it exceeds the 10-second bucket limit (Henning et al., 2024).
+The higher latency is due to the records buffering in Kafka Streams’ repartitioning topic.
+Both frameworks’ executions are similar in the throughput until the first failure occurs.
+After the failure, Flink recovered faster in less than a minute.
+Kafka Streams took longer to reach a higher output throughput that remained unstable, which is due to the lag accumulated and being slowly processed.
+This aligns with Performance observation 2 and Fault recovery observation 2.
+Given that the scenario requiring exactly-once semantics is more critical, Kafka Streams’ suboptimal rebalance strategy has a greater impact.
+7. Fault Recovery Stability
+This section presents the results of fault recovery stability under recurrent failures injected.
+For the sake of visual clarity, note the different ranges and scales on the y-axis.
+For conciseness, we limit the discussion to the results with ALOS guarantees as those are the most complete and insightful.
 Figure 7 complements the results from Fig.
@@
 Fig.
-6(a) shows that despite some fluctuations in the input throughput of Flink’s execution, it achieved a stable behavior under recurrent failures.
+7(a) shows that despite some fluctuations in the input throughput of Flink’s execution, it achieved a stable behavior under recurrent failures.
 Figure 7(b) shows Kafka Streams execution in a scenario where one random worker instance is recurrently killed.
@@
 Moreover, the fluctuations in lag and throughput have a detrimental impact on the latencies, evincing limitations in Kafka Streams’ fault recovery.
-On the one hand, in the recovery of the sixth failure shown in 6(b), the lag accumulated as well as the throughput are lower, and the latency increase is not disruptive with the p50 being kept within the boundaries.
+On the one hand, in the recovery of the sixth failure shown in 7(b), the lag accumulated as well as the throughput are lower, and the latency increase is not disruptive with the p50 being kept within the boundaries.
 On the other hand, following the seventh failure, the performance degradation becomes critical.
 Consequently, the execution takes around 4 minutes to return to a state comparable to the one before the failure.
-In terms of reasons for the observed Kafka Streams’ fault recovery instabilities, the current understanding is that it is mostly due to which worker instance the failure is injected (e.g., if the consumer group leader is affected or not) and how the subsequent(s) repartitioning(s) are enacted.
+In terms of reasons for the observed Kafka Streams’ fault recovery instabilities, the current understanding is that it is mostly due to which worker instance the failure is injected (e.g., if the consumer group leader is affected or not) and how the subsequent rebalances are enacted.
 On the one hand, this outcome is aligned with Fault recovery observation 2.
@@
 8, one can note that Flink and Spark Structured streaming recoveries are similar in all failures.
-Kafka Streams, on the contrary, result in failures having (once again) different impacts on its performance, mostly in terms of latency, lag accumulation, and lag reduction.
+Kafka Streams, on the contrary, result in failures having again different impacts on its performance, mostly in terms of latency, lag accumulation, and lag reduction.
 We can then conclude that Fault recovery observation 2 also applies to worker instances crashes.
-(a) Flink
-(b) Kafka Streams
-(c) Spark Structured Streaming
-Figure 8. ALOS recovery: Two random pods executing worker instances crash for 5 minutes, repeated 5 times every 10 minutes.
 8. Fault Recovery Time
-The precise fault recovery time is another relevant metric that complements the results from Section 6 and Section 7.
-Considering that there is a lack of approaches and tools to automate the measurements of fault recovery times across many failures, we designed a simple detector based on moving average and standard deviation to automatically detect the failure as well as the time to recovery.
-The detector extracts the expected average and standard deviation for both throughput and latency based on the regular processing behavior that follows the warm-up period.
-The time frame ranged between 2 and 5 minutes, and 3 and 6 minutes, for throughput and latency, respectively.
+Estimating fault recovery time is another relevant metric that complements the results from Section 6 and Section 7.
+Considering that there is a lack of approaches and tools to automate the measurements of fault recovery times across multiple failures, we designed a simple detector based on moving average and standard deviation to automatically detect the failure as well as the time to recovery.
+The detector extracts the expected average and standard deviation for the throughput based on the normal processing behavior.
+We define the normal processing period as the interval between the execution warm-up and the injection of the first failure, which can vary depending on the time required for warm-up and the timing of the first failure.
+In our analysis, the normal processing period is between the second and fifth minute after the start of execution across all frameworks.
+In short, the detector marks a point of failure injection when the moving average of a window of 10 data points and its standard deviation are two times above the one from the normal processing period.
+The standard deviation threshold and window size are parametric values defined for all frameworks.
+Afterward, another point of failure was strictly marked every 5 minutes, which is the interval between failures used in our experiments.
+The recovery time was detected when the values returned within a threshold of 15% for a time frame of at least 40 seconds.
+We expect throughput to stabilize after the recovery time, which should be only detected when the execution returns to a normal state.
+It is important to note that our approach for detection failures and their recovery was configured to automate our measurements.
+Considering that it is parametric and customizable, we expect it to be also applicable to wider scenarios.
+[^7] The tooling and all the parameters used are available as supplementary material (Vogel et al., 2024b).
+Moreover, we believe that there are no ground truths about these parameter values used.
+In fact, we focus more on having a formal parametric method to automate the measurements that is less arbitrary and avoids biased analysis.
+The values detected can be an estimation of the recovery time that we expect to be ascertained with visual inspections in the metrics traces.
 Figure 9 exemplifies the trace of automatically detected failures (of 2-pods kill) and their recoveries.
-In short, if the moving average and standard deviation exceeded a given threshold (equal for all frameworks), then the detector would mark it as the point of failure injection.
-Afterward, another point of failure was strictly marked every 5 minutes.
-Similarly, the recovery time was detected when the estimators returned within the same threshold for a larger time frame, since once stability is achieved, throughput and latency should remain constant.
-Considering that the remaining experiments have a high correlation, for conciseness they are not shown here.
-We will provide the additional results and the parameters used for the thresholds in supplemental material.
-Figure 9. Characterization of the automatic fault recovery times detection in the ALOS scenario of 2-pods kill.
-Figure 10 summarizes the results of the fault recovery times.
-Noteworthy, throughput and latency (p90) recovery time are highly correlated.
-Flink achieves the fastest overall recovery time for both metrics.
-An exception is latency on 1-pod kill latency where Spark has a slightly lower median recovery ($m_{flink}=69.5$ seconds, $m_{spark}=68.5$ seconds) but at the cost of higher variability due to the observed spark’s fluctuations in throughput and latency.
-Kafka Streams’ recovery times are longer and with higher variability, consistently with the discussion presented in Section 6.
-There are cases where full recovery was not achieved in the 5 minutes before the next failure was triggered, which is why the median can be high in some Kafka Streams’ results.
-Fault recovery observation 6 is that Flink and Spark Structured Streaming provide faster fault recoveries, and Flink demonstrates the highest level of stability.
-Figure 10. Recovery times.
-9. Concluding Remarks
-Fault tolerance of data-intensive applications is of paramount importance in industry deployments.
+Figure 9.
+Characterization of our approach to automate measurements of fault recovery times in the 2-pods kill ALOS scenario.
+The stable behavior, used as a reference for training, is represented by a black rectangle.
+Failures are highlighted in red, while recovery detection is highlighted in green.
+Figure 10 summarizes the results of the fault recovery times related to the application output throughput.
+Noteworthy, Flink achieves the fastest overall recovery time for both metrics.
+Kafka Streams’ recovery times are longer and with higher variability, which aligns with the discussion provided in Sections 6 and 7.
+In the Kafka Streams scenario of 4 pods killed, a complete recovery was not achieved in the 5-minute window in four out of ten injected failures.
+The same event happened with one Kafka Streams failure in the scenario of 1 pod being killed.
+In such cases, the fault recovery time is defined as the time interval between failures.
+Defining the fault recovery time for these cases as the failure injection interval can impact the estimation because a complete recovery could take longer.
+[^8] The recovery accuracy can be improved with longer intervals between failures, where our follow-up paper provides preliminary results from such a scenario (Vogel et al., 2024a).
+Despite such a scenario, our method captured Kafka Streams’ fault recovery as the slowest due to its limited fault recovery.
+Generally, Fault recovery observation 6 is that Flink and Spark Structured Streaming provide faster fault recoveries, and Flink demonstrates the highest level of stability.
+Considering that fault recovery time is also relevant in terms of the latency metric, we are currently measuring it, and preliminary results are already available in our follow-up paper (Vogel et al., 2024a).
+Figure 10. Output throughput recovery times by framework.
+9. Results Discussion
+Our comprehensive experimental method demonstrates that stream processing frameworks significantly impact application performance and fault recovery.
+Given that fault recovery significantly affects the quality of service, it is a crucial criterion when selecting a stream processing framework for production deployment.
+Section 9.1 shows our key insights from the experiments and Section 9.2 discusses the main findings.
+9.1. Observations Summary and Takeaways
+The following are key insights from our experiments and the observations discussed in Sections 6, 7 and 8:
+  • Flink, Kafka Streams, and Spark Structured Streaming executions are resilient to recover from failures of killed and crashed worker instances, even from failures affecting many worker instances.
+Yet, the impact of their recoveries on the applications’ performance varies in terms of the application’s performance metrics, recovery time, and how stable the recoveries are across recurrent failures.
+  • In use cases where at-least-once guarantees are suitable, Flink and Spark Structured Streaming provide a consistently fast and stable recovery.
+Spark Structured Streaming can be a suitable alternative when low latency is not required.
+On the other hand, Kafka Streams takes longer to recover from failures and has volatile behavior.
+  • Flink is the most suitable framework for use cases requiring exactly-once semantics (EOS) guarantees, providing fast recovery and stable performance.
+Conversely, Kafka Streams’ executions have lower performance when EOS is enabled, and Spark Structured Streaming has limited native support for guaranteeing EOS in executing applications.
+9.2. Findings discussion
+Although many of the insights we discussed regarding how stream processing frameworks recover from failures align with related literature (van Dongen and van den Poel, 2021; Wang et al., 2021, 2022), our experiments reveal new trends and insights that warrant consideration.
+To the best of our knowledge, the following are key novel findings:
+  • $\star$
+    Finding: Kafka Streams provides limited fault recovery.
+  • Rationale: The current Kafka Streams’ partition assignment strategy has not been thoroughly evaluated in related approaches (van Dongen and van den Poel, 2021; Wang et al., 2021).
+Kafka Streams was only tested under failures in (van Dongen and van den Poel, 2021) with a single failure injected.
+Our experimental method reached this finding due to the recurrent injection of different failures and recovery time measurements.
+  • $\star$
+    Finding: Flink currently provides fast and stable recovery.
+  • Rationale: Measuring the stability across recurrent failures is a novelty of our experimental method, which also covers more recent enhancements implemented on Flink.
+Relevant enhancements in Flink include the reactive mode, as well as optimizations in checkpointing and load management (Fragkoulis et al., 2024).
+10. Conclusions
+Ensuring fault tolerance in data-intensive, event-driven applications is crucial for successful industry deployments.
 In this paper, we contribute with a more comprehensive evaluation method and fault recovery analysis.
 Such an evaluation was enabled by ShuffleBench’s (Henning et al., 2024) flexible deployment, parameters, and extension to fault recovery metrics.
-9.1. Key Observations and Takeaways
-We concluded that much has changed compared to the literature on how stream processing frameworks recover from failures.
-This motivates the need for benchmarking studies.
-The following are the key insights from the extensive experimental analysis:
-  • Flink, Kafka Streams, and Spark Structured Streaming are resilient to different types and degrees of failures. However, the frameworks have different impacts on the applications running.
-  • In use cases where at-least-once semantics guarantees are required, Flink provides currently the best recovery, which is both fast and stable.
-Spark Structured Streaming can also provide consistent fault recovery for applications where latency is not a critical requirement.
-Kafka Streams takes longer to recover from failures and presents volatile behavior.
-  • Flink is the best candidate when exactly-once semantics guarantees are necessary.
-On the other hand, Kafka Streams has limited performance, and Spark Structured Streaming has limited native support for EOS.
 10.1. Results Implications
-The fault recovery capabilities should be a criterion for choosing a stream processing framework.
-Therefore, one of the main implications of our findings is that Kafka Streams, despite its growing interest, needs improvements in terms of fault recovery.
-This implication is relevant because our findings indicate that in contrast to what was reported in the past (van Dongen and van den Poel, 2021), Kafka Streams’ unstable recovery can compromise QoS and service level agreements.
-The finding that Kafka Streams faces performance instabilities after failures due to its current repartitioning strategy has relevant implications.
+One of the main implications of our findings is that Kafka Streams, despite its growing interest, needs improvements in terms of fault recovery.
+This implication is relevant because our findings indicate that in contrast to what was reported in the past (van Dongen and van den Poel, 2021), Kafka Streams has unstable recovery that can compromise QoS and service level agreements.
+The finding that Kafka Streams faces performance instabilities after failures due to its current rebalancing strategy has relevant implications.
 We intend to share our findings with Kafka Streams’ community and contribute with future enhancements.
@@
 In fact, our results indicate significant contrasts in recoveries across failures (see Section 8), which can impact the applications’ performance, QoS, and user experience.
-With the significant impact of fault recovery on applications’ performance and QoS, we argue that modern performance benchmarking should be extended to support more realistic scenarios.
-We believe that performance is still a major goal, but it should be continuously improved.
-This would be a starting point to better representing real-world scalable computing deployments in clouds.
-Fault recovery has a high impact on advanced optimizations necessary in modern stream processing, such as run-time self-adaptations for auto-scaling.
-This is because applying the runtime modifications involves recovery.
-Therefore, our findings are also directly applicable to auto-scaling.
-We also believe that our experimental method is generic enough to be useful for evaluating other domains.
-This is mostly because our method is comprehensive and includes insights into how chaos engineering can be applied to long-running computing systems.
+The insights from our experimental results can also impact on advanced optimizations necessary in modern stream processing, such as run-time self-adaptations for auto-scaling.
+This connection arises because implementing run-time adaptations in distributed executions inherently involves fault recovery procedures.
+Considering that our experimental method is comprehensive and includes insights into how chaos engineering can be applied to long-running computing systems, it is expected to be generic enough to be useful for evaluating other domains.
 10.2. Limitations and Threats to Validity
 We expect all threats to the internal validity discussed in our previous works to also apply to our results (Henning et al., 2024).
-In addition to known threats to external validity, the fault measurements have additional configurations and executions threats.
+In addition to known threats to external validity, the fault measurements have additional configurations and execution threats.
 The failure injection interval was defined according to insights from real-world industry deployments, where there can be hundreds or thousands of worker instances, thus increasing failure likelihood.
@@
 10.3. Future Work
-We believe that future analyses to extend ours are still needed.
+We expect that future analyses to extend ours are still needed.
+With the significant impact of fault recovery on applications’ performance and QoS, we argue that modern performance benchmarking should be extended to support more realistic scenarios.
+Performance is still a major goal, but it should be continuously improved.
+This would be a starting point to better representing real-world scalable computing deployments in clouds.
 A relevant lesson learned from the experiments is that 5 minutes between two injected failures is not enough to measure the recovery time in all scenarios.
@@
   • Further analysis of the frameworks’ output correctness under failures with different fault tolerance semantic guarantees.
+  • Extend the measurements to cover execution costs with the stream processing frameworks.
 Acknowledgements.
-We would like to thank Johannes Kepler University Linz and Dynatrace for co-funding this research. We thank Bernhard Kepplinger for providing insights on the method for failure detection.
+We would like to thank Johannes Kepler University Linz and Dynatrace for co-funding this research.
+We thank Bernhard Kepplinger for providing insights into the method for automating the measurement of failure recovery times.